OpenStudy (is3535):

Which circles have an area of 9 m2? Choose all answers that are correct. A. a circle with a radius of 3 m B. a circle with a diameter of 6 m C. a circle with a radius of 9 m D. a circle with a diameter of 9 m Question Resources

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm A}=\pi {\rm r}^2 }\) \(\large\color{black}{ \displaystyle {\rm 9}=\pi {\rm r}^2 }\) \(\large\color{black}{ \displaystyle \frac{{\rm 9}}{\pi}={\rm r}^2 }\) \(\large\color{black}{ \displaystyle \sqrt{\frac{{\rm 9}}{\pi}}={\rm r} }\)

OpenStudy (is3535):

so b and d

OpenStudy (is3535):

@SolomonZelman

OpenStudy (is3535):

@Evil-cow-8

OpenStudy (is3535):

@ElonaSushchik

OpenStudy (is3535):

@ayyookyndall

OpenStudy (is3535):

@isry98